OpenStudy (anonymous):

How do you solve sin(2arccos7/13)

OpenStudy (anonymous):

First, recall the double angle identity for sine: \[\sin\left(2\arccos\frac{7}{13}\right)=2\sin\left(\arccos\frac{7}{13}\right)\cos\left(\arccos\frac{7}{13}\right)\] The strategy is to regard the arccos part as an angle itself: \[\theta=\arccos\frac{7}{13}~~\iff~~\cos\theta=\frac{7}{13}\] This gives you the following type of triangle:  Find the missing side, then you'll have everything you need to find the sine and cosine of \(\theta\).